Define a relation `~' on R by 'a ~ b if a — b is an integer'. Show that this is an equivalence relation. Give the equivalence classes of 0 and √2 .


1
Expert's answer
2020-12-21T17:31:43-0500

Let us show that a relation '"\sim"' on "\mathbb R" by '"a \sim b" if "a - b" is an integer' is an equivalence relation, that is reflexive, symmetric and transitive.


For each "a\in\mathbb R", "a-a=0\in\mathbb Z" , and thus "a\sim a" each "a\in\mathbb R", and the relation is reflexive.


If "a\sim b", then "a-b\in\mathbb Z". It follows that "b-a=-(a-b)\in\mathbb Z", and thus "b\sim a." Consequently, the relation is symmetric.


Let "a\sim b" and "b\sim c". Then "a-b\in\mathbb Z, \ b-c\in\mathbb Z". It follows that "a-c=(a-b)+(b-c)\in\mathbb Z", and thus "a\sim c". Therefore, the relation is transitive.


By defenition, the equivalence class of an element "a" is "[a]=\{b\in\mathbb R\ |\ b-a\in\mathbb Z\}".


Let us give the equivalence classes of 0 and "\sqrt{2}" :


"[0]=\{b\in\mathbb R\ |\ b-0\in\mathbb Z\}=\{b\in\mathbb R\ |\ b\in\mathbb Z\}=\mathbb Z,"


"[\sqrt{2}]=\{b\in\mathbb R\ |\ b-\sqrt{2}\in\mathbb Z\}=\{b\in\mathbb R\ |\ b=\sqrt{2}+n,\ n\in\mathbb Z\}=\sqrt{2}+\mathbb Z."